QUESTION 5

  1. On May 11 of the current year, your calendar year firm purchases a used machine (7-year property) for $10,000; the machine has an estimated salvage value of $1,000. If the machine is the only fixed asset purchased in the current year, and no Section 179 deduction or bonus depreciation is taken, what is your firm’s maximum tax deduction for depreciation this year?

    $9,000

    $1,286

    $1,429

    $10,000

3 points   

QUESTION 6

  1. Generally, under MACRS, the recovery period for office furniture (e.g., desks) is:

    3 years

    7 years

    4 years

    5 years
Answer #1

5. Depreciation = Cost * 14.29% = 10000 * 14.29% = $1429 Option B

6. 7 Years Option B